### v1

# A Comprehensive Evaluation on Event Reasoning of Large Language Models

Abstract
Event reasoning is a fundamental ability that underlies many applications. It requires event schema knowledge to perform global reasoning and needs to deal with the diversity of the inter-event relations and the reasoning paradigms. How well LLMs accomplish event reasoning on various relations and reasoning paradigms remains unknown. To mitigate this disparity, we comprehensively evaluate the abilities of event reasoning of LLMs. We introduce a novel benchmark for EValuation of EVent reasoning. consists of two levels of evaluation of schema and instance and is comprehensive in relations and reasoning paradigms. We conduct extensive experiments on . We find that LLMs have abilities to accomplish event reasoning but their performances are far from satisfactory. We also notice the imbalance of event reasoning abilities in LLMs. Besides, LLMs have event schema knowledge, however, they’re not aligned with humans on how to utilize the knowledge. Based on these findings, we introduce two methods to guide the LLMs to utilize the event schema knowledge. Both methods achieve improvements. Code and Dataset are available on https://github.com/TZWwww/EV2.
A Comprehensive Evaluation on Event Reasoning of Large
Language Models
Zhengwei Tao12 Zhi Jin12††thanks: *Corresponding authors. Yifan Zhang12 Xiancai Chen12 Xiaoying Bai3∗ Yue Fang12 Haiyan Zhao12 Jia Li12 Chongyang Tao4 1Key Laboratory of High Confidence Software Technologies (PKU), MOE, China 2School of Computer Science, Peking University 3Advanced Institute of Big Data 4Beihang University {tttzw, xiancaich, yifanzhang, y.fang}@stu.pku.edu.cn, baixy@aibd.ac.cn {zhijin,zhhy.sei, lijiaa}@pku.edu.cn chongyang@buaa.edu.cn
1 Introduction
Events are instances or occurrences that form the basic semantic building units encompassing the meanings of Activities, Accomplishments, Achievements, and States Vendler (1957). Event Reasoning is the ability to process and analyze events and their complex interconnections. Compared with other abilities, event reasoning is unique in some aspects. Firstly, it requires knowledge in the form of event schemas, capturing the progress of event evolution in scenarios, then performing global reasoning Li et al. (2021a); Mao et al. (2021). As shown in Figure 1, each event instance is associated with an event type. All event types and their relations form the event schema knowledge which reflects the logic and mechanism of event evolution. Knowing “Memory” would often happen after “Learn” can help answer the reasoning question. Second, the inter-event relations and reasoning paradigms are various. Event reasoning incorporates reasoning events according to a certain relation Du et al. (2022); Sap et al. (2019b) and reasoning inter-event relations Ning et al. (2018); Caselli and Vossen (2017). The queried relations are diversified such as causality Roemmele et al. (2011), temporality Zhou et al. (2019), and hierachy Glavaš et al. (2014). There are various paradigms such as reasoning the event or the inter-relation.
As a fundamental competency within LLMs, event reasoning supports a multitude of Natural Language Processing (NLP) tasks, including recommendation engines Yang et al. (2020), interactive question-answer systems Souza Costa et al. (2020), and AI Agents Liu et al. (2023). Therefore, the enhancement of event reasoning abilities is essential for the advancement of LLMs.
LLMs like LLAMA Touvron et al. (2023) series and GPT series Brown et al. (2020) have demonstrated exceptional accomplishments in various natural language reasoning Bang et al. (2023); Xu et al. (2023b). Existing research has evaluated a broad spectrum of reasoning abilities of LLMs such as commonsence Bian et al. (2023), sentence relations Chan et al. (2023), and math Arora et al. (2023). However, studies on the comprehensive evaluation of event reasoning of LLMs are scarce. Current works only focus on instance-level events, resulting in unclearness of how LLMs understand and utilize the event schema knowledge Chan et al. (2023). Besides, they ignore the diversity of relations and paradigms Yuan et al. (2023). These disparities hinge on the development of such crucial abilities of LLMs.
In this paper, we comprehensively evaluate event reasoning in knowledge and abilities. Since there are existing datasets that are comprehensive in relations and paradigms, and can cover both levels of schema and instance, we introduce a novel benchmark for the EValuation of EVent reasoning. is featured in evaluating both aligned schema-level and instance-level. The schema-level evaluation investigates the event schema knowledge of LLMs while the instance-level testifies the event reasoning abilities. Besides, to evaluate event reasoning in various types of relation and reasoning paradigms, includes two event reasoning tasks, namely Contextual Event Classification (CEC) and Contextual Relation Reasoning (CRR) as shown in Figure 1. is constructed from both GPT generation and human annotation. Utilizing , we comprehensively evaluate how well LLMs do event reasoning in terms of abilities and knowledge. Specifically, we mainly explore four research questions: 1) How proficient abilities of event reasoning do LLMs have? 2) To what extent do LLMs have the event schema knowledge? 3) Are LLMs aligned with humans in leveraging event schema knowledge? 4) Can LLMs perform better event reasoning with explicit guidance of leveraging event schema knowledge?
We conduct extensive experiments on to answer these questions. The results provide insights into event reasoning that: 1) LLMs have the abilities of event reasoning, but are far from satisfactory and are imbalanced in different relations and reasoning paradigms. 2) LLMs have event schema knowledge. They can answer the schema-level questions with similar accuracy to the instance-level questions. However, the development of schema-level abilities falls behind those of instance-level. 3) LLMs are not aligned with humans in the aspect of leveraging event schema knowledge. 4) Based on the findings, we design two mentoring methods to guide the LLMs to utilize event schema knowledge. One is to directly add event schema knowledge to the prompt. The second is guiding in a chain-of-thought format. With the designed guidances for utilizing event schema knowledge, LLMs can perform better event reasoning. Especially with direct guidance, LLMs get significant improvements.
We summarize our contributions as follows:
We evaluate event reasoning in both levels of schema and instance, and various relations and paradigms.
We construct a novel benchmark which features two levels of evaluation and comprehensive in relations and reasoning paradigms. We conduct extensive experiments to probe how LLMs perform event reasoning.
We conclude several insights. Based on our findings, we design mentoring methods to guide LLMs to utilize event schema knowledge which achieves improvements in event reasoning.
2 Problem Formulation
Event reasoning is to anticipate the occurrences of certain relations or deduce interrelated correlations Tao et al. (2023a). The relations encompass causality Du et al. (2022), temporality Zhou et al. (2019), and hierarchy Glavaš et al. (2014).
Event reasoning requires comprehension of event schema knowledge. An event schema of a scenario is a schema-level graph 111Superscript represents schema level., where is the set of event types and is the set of relations between events. Each edge in is a relation triplet standing for that there is the relation between and . With instantiation, we have the instance-level event graph 222Superscript represents instance level.. An instance event has an event type but with detailed event arguments and context Mitchell (2005). The nodes and edges of these two graphs are corresponding, namely, each triplet in has a corresponding triplet in with the same inter-relation. In both levels, we consider totally six relation types, namely {Causes, IsResult, Before, After, IsSubevent, HasSubevent}.
consists of two event reasoning paradigms for both levels of schema and instance. The first is Contextual Event Classification (CEC) and the second is Contextualized Relation Reasoning (CRR).
CEC
Given graph , either schema- or instance-level, queried event , and target relation , CEC requires the model to answer an event :
is the model, is the candidate event set. CEC evaluates the model’s comprehension of event semantics and structure.
CRR
Given graph , either schema- or instance-level, two queried events , CRR requires to determine the relation between them:
CRR evaluates the understanding of event relations.
In both schema and instance levels, has CEC and CRR tasks. Schema-level tasks require models to be rich in knowledge while tasks for instance need models to process detailed information.
3 Benchmark Construction
To create the benchmark, we curate a comprehensive dataset through a three-stage process. Initially, the schema graph is established. Then, GPT4 is employed to generate the instance graph . Lastly, human annotators are tasked with creating questions from and .
3.1 Schema Graph Construction
We leverage EECKG Wang et al. (2022b) to ensure a diverse range of event types in our schema. EECKG combines rule-based reasoning with crowdsourced insights, built on ConceptNet’s structure. Nodes in EECKG represent verb phrases as events, and edges denote inter-event relations, focusing on Causes333The direction is that the head event causes the tail event. Other relations are the same., Before, and HasSubevent.
Our objective mandates that the nodes within should represent event types. Therefore, we filter EECKG nodes, removing concrete event instances. Preference is given to nodes with at most two words, as longer descriptions tend to include specific details. For events with fewer than two words, we use GPT4 to enhance our selection, ensuring the appropriate abstraction level for our schema graph with the following prompt:
s
![[Uncaptioned image]](x2.png)
We identify a subset of remaining events that are too generic. To refine the event selection, we also exclude the most frequent events from our subset to avoid generic events.
We then dissect the interconnected EECKG into separate components, each representing a distinct scenario. To prevent semantic drift, we carefully control the size of each component. Starting from a node, we conduct a random walk until the number of nodes surpasses a set threshold, thus defining a component. This process is executed for all nodes to gather all components, as detailed in Algorithm 1. Post-extraction, we eliminate cycles to convert these structures into DAGs.
EECKG only contains forward event evolution relations such as Causes. We further include components of backward relations. We generate a reversed version for each component by inverting edge directions and replacing relations with their opposites: IsResult, After, and IsSubevent. This creates the backward components.
In preparation for constructing tasks for CEC and CRR, we label two events for each component. We sample three event pairs per component with a maximum inter-path length of four, utilizing their predecessors as background events. These pairs and background events form a schema graph. When the path length between and is two, the direct relation serves as the queried relation; for longer paths, we deduce the relation using Table 1. We construct a schema graph, queried event pair, and their relation .
3.2 Instance Graph Construction
We next harvest instance graph for each schema graph . For each node , we ask GPT4 to generate using the following prompt:
We inherit the relations of and obtain . We naturally obtain the instances of and .
3.3 Question Construction
The last step is to construct questions of CEC and CRR in both schema and instance levels. For CEC, regarding schema and instance head events as the query and the tail as an answer, we ask GPT4 to generate 15 possible candidate instance events with their event types.
We then recruit 8 well-educated human annotators. Their missions are:
1)
Revise or discard if not valid. Ensure the events are abstract, the relations are correct, there’s no scenario shifting in .
2)
Revise or discard if not valid. Ensure the events are concrete, the relations are correct, and the whole scenario of is coherent and has no shifting.
3)
Choose three proper negative candidate events with their event types. Ensure answering the question should rely on the context events.
We use the schema part of annotation as the schema-level questions and the instance part as instance-level questions. Then we complete CEC.
For CRR, we regard and as queried events and use the relation between them as the answer to form the schema-level question. For instance part, we adopt a similar way.
Our CEC task is a 4-way multiple-choice task. The CRR is a 3-way multiple-choice task. In CRR, the choices for temporal, causal, and hierarchy relations are [Before, After, Vague], [Causes, IsResult, None], and [IsSubevent, HasSubevent, None] respectively. We show examples of both tasks in Figure 1. We report the number of each task and the average nodes and edges of in Figure 2.
| Rule | Induction |
| (Before)+ | Before |
| (After)+ | After |
| (Before)⋆(Causes)+(Before)⋆ | Before |
| (After)⋆(IsResult)+(After)⋆ | After |
| (Before)⋆(HasSubevent)+( Before)⋆ | Before |
| (Causes)⋆(HasSubevent)+( Causes)⋆ | Causes |
| (After)⋆(IsSubevent)+( After)⋆ | After |
| (IsResult)⋆(IsSubevent)+( IsResult)⋆ | IsResult |
| S-CEC | I-CEC | S-CRR | S-CRR | AVG N | AVG E |
| 492 | 558 | 767 | 835 | 3.62 | 2.78 |
3.4 Quality Inspection
We recruit other human annotators to inspect the quality of . We sample 100 data for all tasks. We ask them to give two scores for each sample:
Correct: Rate 1 if correct, otherwise rate 0.
Contextualized: Rate 1 if the answer relies on the context events, otherwise rate 0.
Finally, we get 91% for Correct and 92% for Contextualized. Human examination testifies that is qualified. Besides, context events count.
3.5 Existing Dataset Comparison
We compare our benchmark to existing related datasets. We show detailed comparison in Table 3. Our benchmark is the only one that is for contextualized event reasoning of various relations and paradigms on both schema and instance levels.
| Dataset | L | C | M-R | M-P |
| ALTLEXHidey (2016) | I | ✗ | ✗ | ✗ |
| ASERZhang et al. (2020) | S | ✗ | ✓ | ✗ |
| ATOMICSap et al. (2019a) | S | ✗ | ✓ | ✗ |
| COPARoemmele et al. (2011) | I | ✗ | ✗ | ✗ |
| CQABondarenko et al. (2022) | I | ✓ | ✓ | ✗ |
| ECAREDu et al. (2022) | I | ✗ | ✗ | ✗ |
| ESLCaselli and Vossen (2017) | I | ✓ | ✗ | ✗ |
| ESTERHan et al. (2021) | I | ✓ | ✓ | ✗ |
| HIEVEGlavaš et al. (2014) | I | ✓ | ✗ | ✗ |
| KAIROSLi et al. (2021a) | S | ✓ | ✗ | ✗ |
| LDC2020E25Li et al. (2021a) | S | ✓ | ✗ | ✗ |
| MATRESNing et al. (2018) | I | ✓ | ✗ | ✗ |
| MAVEN-EREWang et al. (2022a) | I | ✓ | ✓ | ✗ |
| MCNCGranroth-Wilding (2016) | I | ✓ | ✗ | ✗ |
| MCTACOZhou et al. (2019) | I | ✓ | ✗ | ✗ |
| REDO’Gorman et al. (2016) | I | ✓ | ✓ | ✗ |
| SCITELi et al. (2021b) | I | ✓ | ✗ | ✗ |
| SCTMostafazadeh et al. (2016) | I | ✓ | ✗ | ✗ |
| SocialIQASap et al. (2019b) | I | ✓ | ✓ | ✗ |
| TB-DenseCassidy et al. (2014) | I | ✓ | ✗ | ✗ |
| TRACIEZhou et al. (2020) | I | ✓ | ✗ | ✗ |
| S I | ✓ | ✓ | ✓ |
4 Experiments
4.1 Evaluated LLMs
We evaluate 9 LLMs on event reasoning. For the open-source models, we evaluate their chat-version. For the closed-source models, we utilize their official APIs to conduct performance evaluations. Specifically, we employ the gpt-4-0125-preview version as the GPT4 and the gpt-3.5-turbo-1106 version as GPT3.5 in our experiments. For the open-source models, we include Qwen1.5-7B Bai et al. (2023), Mistral-7B Jiang et al. (2023), Baichuan-2-7B Yang et al. (2023), Llama2-7B Touvron et al. (2023), WizardLM-7B Xu et al. (2023a), Vicuna-7B Chiang et al. (2023), and Alpaca-7B Taori et al. (2023). Without loss of generosity, we use the model names to refer to the chat versions in the rest of our paper. For all evaluated LLMs, we use the same prompt. We show prompts in Figure 5-8 in the Appendix.
5 Results and Findings
5.1 How proficient abilities of event reasoning do LLMs have?
In this part, we mainly probe the abilities of how existing LLMs complete the event reasoning of the instance level.
LLMs have the abilities of event reasoning, but even the strongest GPT-4 is far from satisfactory.
We evaluate CEC and CRR at the instance level. We show the results of different relations in Figure 2. For CEC, GPT4 performs the best. Models like Qwen1.5-7B, Mistral-7B, and GPT3.5 are in the second tier. Qwen1.5-7B and Mistral-7B are both better than GPT3.5. Qwen1.5-7B can even excel GPT4 in the temporal and causal relations. The other models such as WizardLM-7B almost fail, obtaining lower than 40% accuracy. For CRR, GPT4 excels all other models as well. However, unlike CEC, there is no obvious difference in the performance of other models for CRR.
We show the average performance of instance-level CEC and CRR in columns I-CEC and I-CRR in Table 4. We only show models that have basic abilities, namely CEC accuracy above 50.00 or CRR above 40.00 in Table 4, while other models may lack analytical significance. Overall, existing LLMs such as GPT4, and Qwen1.5-7B have CECtain event reasoning abilities. However, even the strongest GPT4 can only achieve 63.80 and 61.20 accuracy in each task showing there’s much room for improvements of event reasoning.
The abilities of LLMs to deal with different relations and reasoning paradigms are unbalanced.
Comparing CEC to CRR, as relation-wise results shown in Figure 2 and average performances in columns I-CEC and I-CRR in Table 4, LLMs perform better for CEC than CRR. We compute the average scores of four listed models in Table 4. We find I-CEC is much higher than I-CRR, with 58.91 to 46.18. The results significantly suggest that CRR is harder than CEC. Existing pretraining and SFT datasets may be biased in paradigms.
We then analyze performances on different relations. As shown in Figure 2, LLMs perform best in causality relation. Then, temporal, and hierarchy relations are tied. That further indicates the imbalance training of different relations. Methods and datasets of balanced abilities on relations are needed. Transferring abilities of different relations could also be feasible Tao et al. (2023b).
This is a crucial finding. Chan et al. (2023) conduct causal event classification such as ECARE Du et al. (2022), and relation reasoning such as MATRES Ning et al. (2018). They directly compare these two groups of results and conclude the gaps are merely from differences in relations. However, they ignore the difference in reasoning paradigms. Leveraging , with disentangling relations and formulations, we investigate event reasoning with less bias.
LLMs excel in forward CEC compared with backward.
We calculate the average scores of forward relation (After, IsResult) and backward relation (Before, Causes). The results are shown in columns I-F, and I-B in Table 4. We find that the average of I-F is significantly better than I-B. It also suggests that the training dataset is unbalanced in relations. Less of the training data is used for backward relations, resulting in poorer performances on those. However, backward relations are important in abduction scenarios. Methods should be designed to enhance such abilities.
CEC improves faster than CRR with model development.
We investigate the improvement trends of CEC and CRR. In Figure 3. When models have poor event reasoning abilities, their performances lie around the balanced line showing no significant differences in tasks. With the development, the CEC improves much faster than CRR such models as GPT3.5, Mistral-7B, and Qwen1.5-7B. This investigation appeals to the need for training in comprehensive event reasoning abilities.
| Model | S-F | S-B | I-F | I-B | S-CEC | I-CEC | S-CRR | I-CRR |
| GPT4 | 54.22 | 55.84 | 65.34 | 61.84 | 55.48 | 63.80 | 52.80 | 61.20 |
| GPT3.5 | 47.88 | 50.94 | 55.68 | 49.12 | 49.79 | 50.18 | 45.37 | 39.52 |
| Qwen | 42.96 | 52.45 | 67.05 | 65.37 | 48.98 | 63.98 | 43.00 | 40.00 |
| Mistral | 44.37 | 52.08 | 67.61 | 53.71 | 48.98 | 57.71 | 46.00 | 44.00 |
| AVG | 47.35 | 52.82 | 63.92 | 57.51 | 50.8 | 58.91 | 46.79 | 46.18 |
5.2 To what extent do LLMs have the event schema knowledge?
In the previous section, we acknowledge that LLMs can complete event reasoning to some extent. However, whether they are endowed with event schema knowledge remains unknown. In this part, we mainly explore to what extent LLMs have the event schema knowledge, i.e. of the schema level.
LLMs have event schema knowledge.
We evaluate CEC and CRR on the schema level. The results are shown in Figure 2, and the average scores are reported in Table 4. We find LLMs already have event schema knowledge and can complete both CEC and CRR tasks at the schema level to some extent. However, in Table 4, we observe that S-CEC lags I-CEC, suggesting that LLMs are more adept at reasoning at the instance level.
Event schema knowledge increases falling behind reasoning at the instance level.
We probe how event schema knowledge increases with the development of LLMs. We depict CEC performance comparisons of LLMs on instance- and schema-level in Figure 4. When the models initially can reason about events, they also have event schema knowledge. At this time, models can perform comparatively or even better in schema-level event reasoning. With the development, models perform instance-level reasoning better than schema-level. It indicates that the accumulation of event schema knowledge falls behind the reasoning at the instance level. This finding demonstrates that enhancing event schema knowledge may further improve these abilities to obtain better general LLMs.
5.3 Are LLMs aligned with humans in the aspect of leveraging event schema knowledge
In this section, we investigate how LLMs leverage event schema knowledge to complete event reasoning. We first provide the instance-level question for the models and then ask them to generate the required event schema knowledge to solve the task. Then we evaluate the accuracy of the generated event schema knowledge.
Since we have the ground truth event schema knowledge for each question, the only challenge is to guide the LLMs to generate in a similar format for calculating accuracy. The instruction of our prompt first asks LLMs to generate the event types of each instance event in data. Based on the event types, it requires the LLMs to further generate relation triplets needed for the question.
However, we find the LLMs would generate event types of different words but correct contents. To mitigate this problem, we prepare a list of candidate event types for each data to make it a classification setting. To keep the task difficult, we first conduct KMeans clustering on all event types in our dataset444We use all-mpnet-base-v2 for encoding.. We obtain 1000 clusters. For each data, we assign 20 random candidates in total including the correct ones. The negative event types are chosen from different clusters.
After the generation, we calculate the accuracy of event types and F1-scores of relation triplets respectively comparing with the human-labeled event schema. We regard a correct triplet if all the head and tail event types and the inter-relation align with the human labels. We show detailed examples in Figures 9-10 in Appendix.
The results are in Table 5. We find only GPT4 can generate correct event types while other models all fail. For relation triplet generation, even GPT4 can not output proper event schemas555GPT4 excels other may be attributed to 1) its better alignment. 2) The dataset is originally generated by GPT4.. It significantly suggests that LLMs may not leverage event schema knowledge as humans when solving event reasoning tasks. Alignment of using such knowledge could further improve the performances.
| CEC | CRR | |||
| ET | Rel | ET | Rel | |
| GPT4 | 70.71 | 37.30 | 70.66 | 49.41 |
| GPT3.5 | 13.43 | 15.78 | 18.55 | 21.14 |
| Mistral-7B | 11.15 | 9.00 | 11.88 | 15.15 |
5.4 Can LLMs perform better event reasoning with explicit guidance of leveraging event schema knowledge?
In the previous section, we find LLMs may not leverage event schema knowledge as human does. It raises an interesting question how well LLMs perform if we guide them to explicitly use such knowledge? In this section, we probe this question.
We design two guiding methods:
Direct: Directly add the event type of each instance event into the prompt.
CoT: Guide the LLMs in a CoT-style to 1) generate the event types of each instance event. 2) reason with the event types. This is a more practical method since we would not know the event types in advance in real scenarios.
| CEC | CRR | |||
| w.o.S | w.t.S | w.o.S | w.t.S | |
| GPT4 | 63.80 | 69.89 (6.09) | 61.2 | 63.11 (1.91) |
| GPT3.5 | 50.18 | 60.92 (10.74) | 39.52 | 45.99 (6.47) |
| Mistral-7B | 57.71 | 63.26 (5.55) | 44.00 | 47.07 (3.07) |
| Llama2-7B | 30.29 | 38.17 (7.88) | 34.00 | 43.35 (9.35) |
| WizardLM | 33.69 | 29.93 (3.76) | 37.00 | 44.91 (7.91) |
| Vicuna-7B | 31.18 | 34.41 (3.23) | 42.00 | 42.40 (0.40) |
We show the performances of direct guidance in Table 6. We find incorporating event schema knowledge significantly improves event reasoning. It shows great potential to solve event reasoning with the fusion of event schema knowledge.
We report the results of the CoT guidance in Table 7. We only report results of GPT-4 since we find other models are unable to follow this instruction. We find in CEC, CoT guidance can improve performance. However, the improvement of CoT lags those of Direct, indicating great space for better methods. Developing advanced guidance for all LLMs remains a challenging research problem. We show the example of prompt and GPT4 generation in Figures 11 and 12 in Appendix.
| CEC | w.o.S | w.t.S |
| GPT4 | 63.80 | 67.92 (4.12) |
6 Related Work
Event Reasoning
Du et al. (2022) aims to select the accurate cause or effect event from candidates. Zhou et al. (2019) serves as a dataset for event temporal reasoning. Current works present a scenario of incorporating counterfactual reasoning Qin et al. (2019, 2020). In addition to single-event relation reasoning, existing works also reason events according to diversified event relations Poria et al. (2021); Han et al. (2021); Yang et al. (2022). Tao et al. (2023b) further unifies datasets of several event-inter relations to transfer event relational knowledge to unseen tasks.
Predicting events necessitates the model to anticipate forthcoming occurrences grounded in the present context Zhao (2021). Mostafazadeh et al. (2016) employs a multiple-choice framework to predict future events by encompassing a diverse range of common-sense connections among events. Guan et al. (2019) establish a dataset oriented towards capturing event logic, enabling the generative prediction of future incidents.
Evaluations for LLMs
Evaluating the capacities of LLMs is the foundation of using and improving them. One group of research evaluates the general abilities of LLMs Hendrycks et al. (2020); Zheng et al. (2023); Zhong et al. (2023); Bang et al. (2023) Besides, existing works evaluate LLMs in specific tasks Bang et al. (2023); Bian et al. (2023); Gao et al. (2023); Wei et al. (2023); Li et al. (2024). Related to event reasoning, Yuan et al. (2023) evaluated the ability to solve event relation extraction. Tao et al. (2023a) present the Event Semantic Processing including the event understanding, reasoning, and prediction of event semantics. Chan et al. (2023) investigates relation reasoning between sentences. Compared with them, we are the first to introduce the evaluation for both schema- and instance-level event reasoning. Moreover, we comprehensively evaluate the performances of various relations and reasoning paradigms.
7 Conclusion
In this paper, we evaluate the event reasoning of LLMs. We introduce a novel benchmark which features both levels of schema and instance. It evaluates event schema knowledge and reasoning abilities. Besides, can be used to comprehensively evaluate the event reasoning in various relations and reasoning paradigms. We conduct extensive experiments on . We obtain many insights such as: 1) LLMs have the abilities of event reasoning, but are far from satisfactory and are unbalanced in different relations and reasoning paradigms. 2) LLMs have event schema knowledge. However, with the development of LLMs, this knowledge increases slowly compared with the increase of abilities of event instance reasoning. 3) LLMs are not aligned with human to leaverage event schema knowledge in event reasoning. 4) Based on the findings, we design two methods, namely Direct and CoT, to guide the LLMs to utilize event schema knowledge. With our designed guidances for utilizing event schema knowledge, LLMs can perform better event reasoning.
Limitations
We guide the LLMs to utilize the event schema knowledge in two ways. The Direct effects most. However, the more practical way CoT falls behind Direct indicating the potential of a better method of guidance. We leave it to future work.
Appendix